OpenStudy (shaleiah):

Question open for 27 mins.

OpenStudy (shaleiah):

How do I find the vertex and axis of symmetry of: y=-2x^2+28x-18

OpenStudy (shaleiah):

@mathstudent55

OpenStudy (phi):

two ways. 1) "complete the square" so that you can write the equation in vertex form y = a( x-h)^2 +k x=h is the axis of symmetry 2) use x= - b/(2a) where a, b, (and c but we don't use it) are from matching your equation to y = a x^2 + b x + c

OpenStudy (michele_laino):

hint: the x-coordinate of the symmetry axis, is: \[\Large {x_V} = - \frac{b}{{2a}} = - \frac{{28}}{{2 \times \left( { - 12} \right)}} = ...?\]

OpenStudy (michele_laino):

more precisely, it is the equation of the symmetry axis and the x-.coordinate of the vertex of the parabola

OpenStudy (shaleiah):

-7/6

OpenStudy (michele_laino):

I got \(7/6\)

OpenStudy (michele_laino):

so, the equation of the symmetry axis, is: \[\Large x = \frac{7}{6}\] and the x-coordinate of the vertex of the parabola, is: \[\Large {x_V} = \frac{7}{6}\]

OpenStudy (michele_laino):

now, in order to get the y-coordinate of the vertex of parabola, we have to replace \(x=7/6\) into the equation of parabola

OpenStudy (michele_laino):

so, we have to compute this: \[\Large \begin{gathered} {y_V} = - 2{x^2} + 28x + 18 = \hfill \\ \hfill \\ = - 2 \cdot {\left( {\frac{7}{6}} \right)^2} + 28 \cdot \left( {\frac{7}{6}} \right) + 18 = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. it is -18: \[\Large \begin{gathered} {y_V} = - 2{x^2} + 28x - 18 = \hfill \\ \hfill \\ = - 2 \cdot {\left( {\frac{7}{6}} \right)^2} + 28 \cdot \left( {\frac{7}{6}} \right) - 18 = ...? \hfill \\ \end{gathered} \]

OpenStudy (shaleiah):

\[\frac{ 215 }{ 18 }\]

OpenStudy (michele_laino):

that's right! so the vertex is the subsequent point: \[\Large V = \left( {\frac{7}{6},\frac{{215}}{{18}}} \right)\]

OpenStudy (michele_laino):

I'm very sorry, I have made a typo, here is the x-coordinate of the vertex: \[\Large {x_V} = - \frac{b}{{2a}} = - \frac{{28}}{{2 \times \left( { - 2} \right)}} = 7\]

OpenStudy (michele_laino):

since the first coefficient it is \(-2\) and not \(-12\)

OpenStudy (michele_laino):

so the y-coordinate of the vertex, is: \[\Large \begin{gathered} {y_V} = - 2{x^2} + 28x - 18 = \hfill \\ \hfill \\ = - 2 \cdot {\left( 7 \right)^2} + 28 \cdot \left( 7 \right) - 18 = 80 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

namely: \[\huge V = \left( {7,80} \right)\]

OpenStudy (shaleiah):

Correct :)

OpenStudy (michele_laino):

:) and the equation of the axis of symmetry, is, of course: \(\huge x=7\)